OpenStudy (anonymous):

-2y - 7 > 11 Part 1: Solve the inequality above. Part 2: Describe the graph of the solution.

OpenStudy (anonymous):

To solve an inequality, pretend it is a normal equation you are trying to solve,

OpenStudy (anonymous):

-2y>18

OpenStudy (anonymous):

Do you follow?@andrewj43

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

Ok, what confused you?

OpenStudy (anonymous):

everything really

OpenStudy (anonymous):

Ok, I will show you again. To solve these, replace the > sign with an equal sign=. So, you would have -2y-7=11. Now, -2y=11+7. -2y=18

OpenStudy (anonymous):

Do you follow so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

We will now divide both sides by 2, so, -y=9

OpenStudy (anonymous):

This is an important step, it is crucial to solving these questions, I will show you now.

OpenStudy (anonymous):

Are you following?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok,